OpenStudy (anonymous):

what is the next term in the geometric sequence 6, -4, 8/3, -16/9

OpenStudy (campbell_st):

compare the terms to find the common ratio does \[\frac{-4}{6} = \frac{-4}{\frac{8}{3}}\] if it does then you have the common ratio... r so the next term will be \[\frac{-16}{3} \times r\] hope it helps

OpenStudy (anonymous):

how did you find that i mean did you solve the ratio?

OpenStudy (anonymous):

what is r

OpenStudy (campbell_st):

just see if both simplify to the same value oops and the 2nd ratio should be \[\frac{\frac{8}{3}}{-4}\]

OpenStudy (campbell_st):

r is the common ratio.... by comparing the ratios -4/6 and (8/3)/-4 you can find what each term is being multiplied by to get the next term

OpenStudy (anonymous):

Oh ok I get it thank you.

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{8}{3}}{-4}\implies \cfrac{\frac{8}{3}}{-\frac{4}{1}}\implies \cfrac{8}{3}\cdot -\cfrac{1}{4}\)

OpenStudy (anonymous):

So what would the answer be because i tried it and got -16=16

OpenStudy (campbell_st):

nope.... it doesn't fit the pattern... so start by simplifying -4/6 what do you get..?

OpenStudy (anonymous):

-2/3

OpenStudy (campbell_st):

great so then check the 3rd term is 3rd term = 2 term times -2/3 which means is \[\frac{8}{3} = -4 \times -\frac{-2}{3}\]

OpenStudy (campbell_st):

if it is then you've found the common ratio...

OpenStudy (anonymous):

my options are A.-8/3 B.-32/27 C.32/27 D.8/3 E.106/9

OpenStudy (jdoe0001):

\(\large \begin{array}{ccccc} 6&-4&\frac{8}{3}&-\frac{16}{9}\\ \hline\\ &\qquad 6\cdot {\color{red}{ -\frac{2}{3}}}&\qquad -4\cdot {\color{red}{ -\frac{2}{3}}}&\qquad \frac{8}{3}\cdot {\color{red}{ -\frac{2}{3}}} \end{array}\)

OpenStudy (jdoe0001):

so... what do you think is next?

OpenStudy (anonymous):

B.-32/27 C.32/27 right?

OpenStudy (jdoe0001):

well. what did you get?

OpenStudy (anonymous):

32/27

OpenStudy (anonymous):

that's what i got

OpenStudy (jdoe0001):

the idea of a sequence is that, the NEXT term, depends on whatever the PREVIOUS term is so you grab the PREVIOUS term and multiply it by some value, that is the "common ratio"

OpenStudy (anonymous):

Thank you

OpenStudy (jdoe0001):

\(\large \begin{array}{ccccc} 6&-4&\frac{8}{3}&-\frac{16}{9}&{\color{blue}{ \frac{32}{27}}}\\ \hline\\ &\qquad 6\cdot {\color{red}{ -\frac{2}{3}}}&\qquad -4\cdot {\color{red}{ -\frac{2}{3}}}&\qquad \frac{8}{3}\cdot {\color{red}{ -\frac{2}{3}}}&\qquad -\frac{16}{9}\cdot {\color{red}{ -\frac{2}{3}}} \end{array}\) 32/27 \(\large \checkmark\)

OpenStudy (anonymous):

There we go thank you again

OpenStudy (jdoe0001):

yw